OpenStudy (anonymous):

If f(x)=4x-7and g(x)=2x+4 evaluate f(x) +g(x) for x =-3 A -52 b-21 c-22 d-15

OpenStudy (sleepyhead314):

so first do f(x) + g(x) plugging in f(x)=4x-7 and g(x)=2x+4 for each so (4x - 7) + (2x + 4) = ? then once you get that ^ plug in (-3) for x :)